OpenStudy (anonymous):

5^x=3^x+2 solve equation express irrational solutions in exact form

OpenStudy (radar):

x=1

OpenStudy (radar):

From inspection. Now how to solve step by step!

OpenStudy (anonymous):

i need to solve step by step that is were i am confused

OpenStudy (radar):

Using logs would end up with decimal values, and the want "exact form"

OpenStudy (radar):

@tiaph, any ideas?

OpenStudy (radar):

It is above my paygrade cassyynn. Maybe repost it and see if the big dogs will bite it lol.

OpenStudy (anonymous):

thanks for trying

OpenStudy (radar):

Sorry I did not solve it. I am going to have to review exponential equations!

OpenStudy (anonymous):

can be written as 2=5^x-3^x. 5^x-3^x... same exponents... so reduce to 2=2^x. x=1 because any other real value of x would make the equation not solvable.

OpenStudy (anonymous):

but if you put 1 back in for 2 it doesn't make a real answer

OpenStudy (radar):

But you don't do that....you only substitute the 1 for x\[5^{1}=3^{1}+2\] 5=5

OpenStudy (radar):

@mariamjones, are you sure:\[5^{x}-3^{x}=2\] Lets say would \[2^{x}+4^{x}=72\] \[6^{x}=72\] I don't think so. In this case x=3. Think about this.